OpenStudy (anonymous):

find the exact value of sin(cos(-13/5)) explain please! i have the answer from the back of the book, which is 4/5 , but i dont know how to get there. Thanks!!

OpenStudy (mathstudent55):

Is it cos(13/5) or arccos(-13/5) ?

OpenStudy (anonymous):

just cos

OpenStudy (anonymous):

sin(cos(-13/5)

OpenStudy (mathstudent55):

It's not this? \(\sin (\cos^{-1} \left(-\dfrac{13}{5} \right))\)

OpenStudy (anonymous):

no

OpenStudy (phi):

how about \[ \sin \left(\cos^{-1}\left(\frac{3}{5}\right) \right) \]

OpenStudy (anonymous):

\[\sin(\cos \frac{ -13 }{ 5 }\]

OpenStudy (anonymous):

can i change it to that... pi?

OpenStudy (mathstudent55):

Is there an "exponent" of -1 on the cosine?

OpenStudy (phi):

the little -1 next to the cos means inverse cosine it means find the angle whose cosine is 3/5

OpenStudy (phi):

it is "special terminology" because it is confusing this same problem might also be written as \[ \sin \left( acos\left( \frac{3}{5}\right) \right) \]

OpenStudy (phi):

or even \[ \sin \left( \arccos\left( \frac{3}{5}\right) \right) \]

OpenStudy (anonymous):

ok that makes more sense.. hold on while i try to figure it out

OpenStudy (anonymous):

so would the 3 be the adjacent side and 5 be the hyp.

OpenStudy (phi):

yes. use pythagoras to find the other leg (the opposite side) the sin of the angle is opp/hyp

OpenStudy (anonymous):

AWESOM!!!!! thank you so much

OpenStudy (anonymous):

can i ask u another question

OpenStudy (phi):

please make it a new post.